OpenStudy (anonymous):

Joe is trying to soup up his dragster. He knows that the time needed for the car to go from 0 to 100 miles per hour varies inversely with the car's horsepower. At 200 horsepower, the car can go from 0 to 100 mph in 12 seconds. How long should the car take if he can increase the horsepower to 240 hp?

ganeshie8 (ganeshie8):

here, pickup time is inversely proportional to horsepower. \[\huge{time\ \alpha\ \frac{1}{horsepower}}\] \[\huge{time\ = \ k\frac{1}{horsepower}}\]

ganeshie8 (ganeshie8):

its given that car takes 12 seconds with 200 horsepower \[\huge{12\ = \ k\frac{1}{200}} \] \[\huge{k = 2400}\]

ganeshie8 (ganeshie8):

plugin "k" value in above formula \[\huge{time\ = \ 2400\frac{1}{horsepower}} \]

ganeshie8 (ganeshie8):

we need to find how much time it takes for the car to pickup 100 miles with a horsepower of 240 just plugin "horsepower = 240" in above formula and simplify

OpenStudy (anonymous):

????????????